# Properties preserved by classes of Chu transforms

### v1

@@ 6. Graphs as Chu spaces
A graph is a pair $\langle V,E\rangle$ such that $E\subseteq{[V]}^{2}$. In this section, we briefly consider two possible realizations of graphs as Chu spaces. At a first glance, a graph could be defined as a Chu space $\langle X,r,X\rangle$ which is simple and undirected, in the sense that:
  1. (1)
    for all $x\in X$, $\langle x,x\rangle\notin r$;
  2. (2)
    for all $x_{0},x_{1}\in X$, if $\langle x_{0},x_{1}\rangle\in r$ then $\langle x_{1},x_{0}\rangle\in r$.
If $(f,g)$ is a Chu transform from a graph $\langle X,r,X\rangle$ to a graph $\langle Y,s,Y\rangle$, then condition (2) implies that $(g,f)$ is also a Chu transform from $\langle Y,s,Y\rangle$ to $\langle X,r,X\rangle$. On the other hand, it is easy to see that, given two graphs $\langle X,r,X\rangle$ and $\langle Y,s,Y\rangle$ which both have an isolated vertex, there always exists a Chu transform from $\langle X,r,X\rangle$ to $\langle Y,s,Y\rangle$.
This observation shows that in this coding the language of flow formulas is rather too weak to express graph properties. In fact, even though the two sorts in the flow formulas are both the same, namely the vertices of the graph, it does not mean we have full access to quantification over them. For instance, once a vertex is quantified as $\exists v^{\mathsf{p}}$ inside a formula, it can only be joined to a vertex of the form $v^{\mathsf{s}}$. One could say that any formula which includes $\exists v^{\mathsf{p}}$ defines a bipartition of the graph, depending on the formula.
For this reason, we are motivated to consider a different coding. The idea is that the points of the Chu space should be the vertices of the graph, whereas the states should consist of all the edges of the graph, plus the empty state. That is, a graph $\langle V,E\rangle$ is represented as a Chu space $\langle V,\in,E\cup\{\emptyset\}\rangle$. To describe Chu transforms in this context, we recall a definition from a recent paper by Stuart Margolis and John Rhodes [21], where strict continuous functions on graphs are introduced.
@@
Lemma 6.2 (Margolis and Rhodes [21, Lemma 2.9]).
@@
(eq-7) $v\in f^{-1}[e^{\prime}]\iff f(v)\in e^{\prime}\iff v\in g(e^{\prime})$
@@
Conversely, if $f^{-1}[e^{\prime}]=g(e^{\prime})$ for each $e^{\prime}\in E^{\prime}\cup\{\emptyset\}$, then again the two equivalences of (eq-7) show that $(f,g)$ is a Chu transform. ∎
@@
From Proposition 6.3 and Remark 6.4, we see that both Chu transforms and dense Chu transforms in this coding give rise to purely graph-theoretic notions. A characterization of the graph properties preserved by such function then follows from Theorem 3.6 and Theorem 4.12, respectively.
@@ 7. Conclusion and further directions
We have investigated the concept of Chu spaces and transforms, coming from category theory, for which we showed the use in various parts of logic and mathematics in general. These include abstract logics, topology (in particular regarding compactness properties, see §4), the theory of ultrafilters and reduced products (§5), and graph theory (§6). We have shown that many classically considered notions in mathematics, such as the Rudin-Keisler order between ultrafilters, are special cases of Chu transforms.
As a main contribution of the paper, in §3 and §4 we have established a number of preservation results for classes of ,$\mathcal{L}^{\mathrm{II}}_{\infty,\infty}$-formulas under various kinds of Chu transforms. Our main theorem in this respect is the characterization of $\mathcal{L}^{\mathrm{II}}_{\omega,\omega}$-formulas preserved under dense Chu transforms, Theorem 4.12.
In future work, we intend to give syntactic characterizations and preservation theorems for two other related category-theoretic concepts: generalized Galois-Tukey connections and Dialectica morphisms. Establishing a more general preservation theorem, along the lines of Keisler [17, Theorem 3], we may treat all such cases in a unified way.
Feferman observed in [13, Remark 2] that his proof of van Benthem’s Theorem 3.6 works also for countable admissible fragments of $L_{\omega_{1},\omega}$. We do not know to what extent the result extends to larger infinitary languages. It is natural to start by looking for counterexamples, so by the following question:
@@
If an ,$\mathcal{L}^{\mathrm{II}}_{\infty,\omega}$-formula is preserved under Chu transforms, must it be logically equivalent to a flow formula?
Another interesting direction is to extend our analysis from $2$-valued Chu spaces to $2^{n}$-valued Chu spaces, namely triples $\langle X,r,A\rangle$ with $r\colon X\times A\to 2^{n}$. The interest in such spaces is motivated by Vaughan Pratt [24, Theorem 14], who constructed a full concrete embedding from the category of $n$-ary relational structures and homomorphisms into the category of $2^{n}$-valued Chu spaces and Chu transforms. The $2^{n}$-valued Chu spaces are of particular interest when studying graphs, since $4$-valued Chu spaces encode graphs much better than the two codings we considered in §6. In particular, it is possible for such a coding to express the notion of a graph homomorphism. Since we have an ambitious hope to formulate and explain Rossman’s homomorphism preservation theorem [25, Theorem 5.16] in our framework, the following question is the natural starting point:
@@
Which properties are preserved by Chu transforms between $2^{n}$-valued Chu spaces?

### v3

@@ 6. Graphs as Chu spaces
In this section, we briefly consider two possible realizations of graphs and related structures as Chu spaces.
6.1. Strictly continuous functions between graphs
A graph is a pair $\langle V,E\rangle$ such that $E\subseteq{[V]}^{2}$. Each graph $\langle V,E\rangle$ can be naturally represented as an extensional Chu space $\langle V,\in,E\cup\{\emptyset\}\rangle$, where the points are the vertices of the graph, the relation is the membership relation, and the states consist of all the edges of the graph plus the empty state. To describe Chu transforms in this context, we recall a definition from a recent paper by Stuart Margolis and John Rhodes [22], where strictly continuous functions on graphs are introduced.
@@ 6.1. Strictly continuous functions between graphs
Lemma 6.2 (Margolis and Rhodes [22, Lemma 2.9]).
@@
(6) $v\in f^{-1}[e^{\prime}]\iff f(v)\in e^{\prime}\iff v\in g(e^{\prime})$
@@
Conversely, if $f^{-1}[e^{\prime}]=g(e^{\prime})$ for each $e^{\prime}\in E^{\prime}\cup\{\emptyset\}$, then again the two equivalences of (6) show that $(f,g)$ is a Chu transform. ∎
@@
From Proposition 6.3 and Remark 6.4, we see that both Chu transforms and dense Chu transforms in this coding give rise to purely graph-theoretic notions.
6.2. Adjointable maps between orthosets with 0
In recent work, Jan Paseka and Thomas Vetterlein [27] introduced orthosets with 0 as a framework to investigate the concept of orthogonality in Hilbert spaces.
Definition 6.5 (Paseka and Vetterlein [27, Definition 2.1]).
An orthoset with 0 is a triple $\langle X,\perp,0\rangle$ where $X$ is a non-empty set, $\perp$ is a binary relation on $X$, and $0\in X$, such that the following conditions are satisfied:
  • for all $x_{0},x_{1}\in X$, if $x_{0}\perp x_{1}$ then $x_{1}\perp x_{0}$;
  • for all $x\in X$, if $x\perp x$ then $x=0$;
  • for all $x\in X$, $0\perp x$.
Remark 6.6.
It is clear from the definition that if $\langle X,\perp,0\rangle$ is an orthoset with 0 then $X\setminus\{0\}$ inherits the structure of a graph, in which two vertices $x_{0}$ and $x_{1}$ are joined by an edge if and only if $x_{0}\perp x_{1}$.
We represent orthosets with 0 as Chu spaces $\langle X,\perp,X\rangle$. In this realization, many properties of an orthoset with 0 can be recast as properties of the associated Chu space. For example, $\langle X,\perp,0\rangle$ is irredundant in the sense of [27, Definition 2.6.(i)] if and only if $\langle X,\perp,X\rangle$ is separated if and only if $\langle X,\perp,X\rangle$ is extensional.
Definition 6.7 (Paseka and Vetterlein [27, Definition 3.1]).
Let $\langle X,\perp,0\rangle$ and $\langle Y,{\perp},0\rangle$ be orthosets with 0. Given $f\colon X\to Y$ and $g\colon Y\to X$, we say that $g$ is an adjoint of $f$ if for every $x\in X$ and every $y\in Y$
$x\perp g(y)\iff f(x)\perp y.$
The following characterization is immediate from the definitions.
Proposition 6.8.
Let $\langle X,\perp,0\rangle$ and $\langle Y,\perp,0\rangle$ be orthosets with 0. For a pair of functions $f\colon X\to Y$ and $g\colon Y\to X$, the following conditions are equivalent:
  • $(f,g)$ is a Chu transform from $\langle X,\perp,X\rangle$ to $\langle Y,\perp,Y\rangle$;
  • $g$ is an adjoint of $f$.
We note that, unlike the cases considered previously, the two conditions of Proposition 6.8 are symmetric in $f$ and $g$, meaning that $(f,g)$ is a Chu transform from $\langle X,\perp,X\rangle$ to $\langle Y,\perp,Y\rangle$ if and only if $(g,f)$ is a Chu transform from $\langle Y,\perp,Y\rangle$ to $\langle X,\perp,X\rangle$.
@@ 7. Conclusion and further directions
We have investigated the concepts of Chu spaces and Chu transforms, coming from category theory, for which we showed the use in various parts of logic and mathematics in general. These include abstract logics, topology (in particular regarding compactness properties, see §4), the theory of ultrafilters (§5), and graph theory (§6). We have shown that many classically considered notions in mathematics, such as the Rudin-Keisler order between ultrafilters, are special cases of Chu transforms.
As a main contribution of the paper, in §3 and §4 we have established a number of preservation results for classes of $\mathcal{L}^{\mathrm{II}}_{\infty,\infty}$-formulas under various kinds of Chu transforms. Our main theorem in this respect is the characterization of $\mathcal{L}^{\mathrm{II}}_{\omega,\omega}$-formulas preserved under dense Chu transforms, Theorem 4.14.
Feferman observed in [15, Remark 2] that his proof of van Benthem’s Theorem 3.6 works also for countable admissible fragments of $\mathcal{L}_{\omega_{1},\omega}$. We do not know to what extent Theorem 4.14 extends to larger infinitary languages.
@@
If an $\mathcal{L}^{\mathrm{II}}_{\infty,\omega}$-formula is preserved under dense Chu transforms, must it be logically equivalent to an inconsistency-flow formula?
To address this question, it is natural to look at the techniques of Barwise and van Benthem [6], which yield a generalized interpolation theorem for $\mathcal{L}_{\infty,\omega}$.
Another interesting direction is to extend our analysis from $2$-valued Chu spaces to $k$-valued Chu spaces, namely triples $\langle X,r,A\rangle$ with $r\colon X\times A\to k$. The interest in such spaces is motivated by Vaughan Pratt [28, Theorem 14], who constructed a full concrete embedding from the category of $n$-ary relational structures and homomorphisms into the category of $2^{n}$-valued Chu spaces and Chu transforms. For this reason, $2^{n}$-valued Chu spaces seem well suited to express the notion of homomorphism between relational structures and the corresponding preservation results, such as Rossman’s homomorphism preservation theorem [30, Theorem 5.16]. More recently, Shengyang Zhong [39] used $3$-valued Chu spaces to formalize the intuition behind quantum logics and Holliday’s fundamental logic [18].
@@
Which properties are preserved by Chu transforms between $k$-valued Chu spaces?
In future work, we also hope to give syntactic characterizations and preservation theorems for two other related category-theoretic concepts: Dialectica morphisms and generalized Galois-Tukey connections.
Acknowledgements
We thank Tom Benhamou for pointing out the connection between Rudin [32, Theorem 1.5] and our Theorem 5.4. We also thank the anonymous referee, whose report was particularly helpful. It is thanks to a suggestion in the report that we considered recursively saturated model pairs and were able to significantly shorten the proof of Theorem 4.14, which originally was proved using the method of elementary chains.